I am trying to solve one last word problem:
Nutrition: One eight-ounce glass of apple juice and one eight-ounce glass of orange juice contain a total of 185 milligrams of vitamin C. Two eight-ounce glasses of apple juice and three eight-ounce glasses of orange juice contain a total of 452 milligrams of vitamin C. How much vitamin C is in an eight-ounce glass of each type of juice?
Thanks so much!!
Denise

Answer by lwsshak3(11628)

let x = number of miligrams of vitamin C in an eight-ounce glass of orange juice
then, 185-x = number of miligrams of vitamin C in an eight-ounce glass of apple juice
two eight-ounce glasses of orange juice and three eight-ounce glasss of apple juice contain 452 milligrams of vitamin C
2x+3(185-x)=452
2x+555-3x=452
-x=452-555
x=103 miligrams of vitamin C
185-x = 82 miligrams of vitamin C
ans: amount of vitamin C in an eight-ounce glass of orange juice is 103 miligrams
amount of vitamin C in an eight-ounce glass of apple juice is 82 miligrams